OpenStudy (vickyisthesmartone123):

What is the value of the ratio of the surface area to the volume of the right prism? A. 1.4 B. 1.07 C. 0.94 D. 0.7

OpenStudy (vickyisthesmartone123):

@amistre64

OpenStudy (amistre64):

you have not provided any prism to determine the results with.

OpenStudy (vickyisthesmartone123):

prooe your answer

OpenStudy (vickyisthesmartone123):

think it c

OpenStudy (amistre64):

SA: V or written another way: SA/V surface area is: basearea + perimeter of base*height volume is: lngth * width * height im not getting C or D

OpenStudy (vickyisthesmartone123):

now that I did that im getting....b

OpenStudy (amistre64):

lets chk: Surface area (6+6+3+3)*5 + 2(6)(3) (18)5 + 2(18) 18(7) = 70+56 = 126 ------------------------ vol 6(3)(5) = 30(3) = 90 ----------------------- 126/90 is not 1.07 either

OpenStudy (vickyisthesmartone123):

well.....................................................................................I got AAAAAAAA it is AAAAA A

OpenStudy (amistre64):

seems like a reasonable pick to me